OpenStudy (anonymous):

WILL FAN AND MEDAL A recipe that makes 6 servings calls for 1/3 cup sugar. Arlene modifies the recipe so that it can serve 9. How many cups of sugar does Arlene need?

OpenStudy (studygurl14):

\[(\frac{1}{3} \div 6)9 = (\frac{1}{3} \times \frac{1}{6})9=\frac{1}{18}\times9=\frac{9}{18}=\frac{1}{2}\]

OpenStudy (anonymous):

\(\sf \Huge 9\ serving (\frac{1}{6\ serving})(\frac{1}{3} cup)=?\)

OpenStudy (anonymous):

thanks @StudyGurl14 ...... again

OpenStudy (studygurl14):

no prob.